# Automted Workload Mapping with 3D Heterogeneous Electronic-Photonic Processing-in-Memory Platform HPiMAP: A Framework for Mapping Hybrid DNNs onto Heterogeneous 3D-Stacked Electronic-Photonic Processing-in-Memory Architectures HPiMAP: A Multi-Objective DNN Mapping Framework on Heterogeneous Electronic-Photonic Processing-in-Memory Architectures

### v1

@@ Abstract
The future of artificial intelligence (AI) acceleration demands a paradigm shift beyond the limitations of purely electronic or photonic architectures. Photonic analog computing delivers unmatched speed and parallelism but struggles with data movement, robustness, and precision. Electronic processing-in-memory (PIM) enables energy-efficient computing by co-locating storage and computation but suffers from endurance and reconfiguration constraints, limiting it to static weight mapping. Neither approach alone achieves the balance needed for adaptive, efficient AI. To break this impasse, we study a hybrid electronic-photonic-PIM computing architecture and introduce H3PIMAP, a heterogeneity-aware mapping framework that seamlessly orchestrates workloads across electronic and optical tiers. By optimizing workload partitioning through a two-stage multi-objective exploration method, H3PIMAP harnesses light speed for high-throughput operations and PIM efficiency for memory-bound tasks. System-level evaluations on language and vision models show H3PIMAP achieves a 2.74 energy efficiency improvement and a 3.47 latency reduction compared to homogeneous architectures and naïve mapping strategies. This proposed framework lays the foundation for hybrid AI accelerators, bridging the gap between electronic and photonic computation for next-generation efficiency and scalability.
I Introduction
@@
Artificial Intelligence (AI), powered by deep neural networks (DNNs), has become integral to a wide array of application domains, e.g., computer vision, natural language processing (NLP), medical diagnostics, and time-series data analytics. These increasingly sophisticated workloads require low latency, high throughput, energy efficiency, and flexibility, objectives that traditional CPUs and GPUs often struggle to meet. To address issues in the notorious "memory wall" and growing computational inefficiency, researchers have explored various domain-specific accelerators that depart from the conventional von Neumann paradigm.
Among many emerging solutions, Processing-in-Memory (PIM) and photonic accelerator systems stand out. PIM harnesses the inherent capability of performing computations directly within or near the memory fabric [1], thereby circumventing the costly data transfers across separate processing and memory units. While PIM can substantially reduce latency and energy consumption, different memory technologies (e.g., SRAM [2], FeFET [3], ReRAM [4], MRAM [5]) each impose unique trade-offs involving footprint, endurance, write latency, and power dissipation. No single technology uniformly excels in all these dimensions, so effective PIM designs require careful technology and architectural selections.
Photonic accelerators, on the other hand, exploit the massive parallelism and high bandwidth of optical signals to process matrix-vector multiplications at unprecedented speeds. By encoding information in light, photonic computing can execute operations such as convolutions or fully connected layers with high throughput. Nonetheless, challenges like thermal crosstalk, process variations, limited data precision, and the overheads of optoelectronic conversion constrain the practical deployment of purely photonic systems. Developing robust calibration and error-compensation strategies remains a critical step toward unlocking the full potential of photonic accelerators.
Another motivation for embracing heterogeneous accelerators stems from the intrinsic heterogeneity in modern DNN workloads, for example, different layer sizes, synaptic sensitivity, input operand properties (dynamic or static), etc. These heterogeneous characteristics influence the suitability and efficiency of individual workload components (e.g., layers/channels) for mapping onto specific processing platforms. For example, mapping dynamic tensor products (attention layers in Transformers) to non-volatile PIM hardware [6] will cause significant endurance concerns. Fully leveraging the potential of heterogeneous accelerator platforms, therefore, demands an automated heterogeneity-aware workload mapping framework toward efficiency-optimal cooperation across hybrid AI accelerators.
In this work, we propose, for the first time, a heterogeneity-aware multi-objective mapping framework to intelligently distribute hybrid DNN workloads across heterogeneous electronic-photonic AI accelerators. Our two-stage mapping flow integrates state-of-the-art architecture and interconnect simulation to explore Pareto-optimal mapping solutions in energy-latency space. Then, our sensitivity-aware workload remapping stage rapidly boosts the mapping’s robustness against various hardware non-ideality to meet the accuracy target. Our key contributions are as follows:
@@
Electronic-Photonic-PIM Accelerator Evaluation Infrastructure: We create the first system modeling and optimization infrastructure integrating SoTA simulators for PIM, photonic accelerators, and network-on-chips, enabling automated design space exploration.
@@
Heterogeneity-aware Mapping Framework: We introduce a two-stage multi-objective mapping framework (H3PIMAP) that considers workload variability and hardware heterogeneity, ensuring rapid mapping space exploration for optimal resource utilization, system performance, and inference accuracy.
Comprehensive Performance Evaluation: H3PIMAP shows 2.74 and 3.47 lower energy and latency over homogeneous systems and naive mapping strategy via evaluation over various NLP and Vision models/datasets.
II Related Work
We discuss relevant work that has been done within the scope of PIM and photonic DNN accelerator architectures. Specifically, we focus on homogeneous architectures solely based on either SRAM, ReRAM or photonic devices, reviewing their advantages and limitations. Table I compares the characteristics of SRAM, ReRAM, and photonic devices. Similar to our baseline ISAAC [7] and TeMPO [8] architectures, we discuss homogeneous PIM and photonic architectures and review prior work done on heterogeneous systems for DNN accelerators.
II-A PIM DNN Accelerators
PIM-based architectures offer an efficient solution for DNN inference by performing computations within the memory, thereby reducing memory traffic and data movement. DNN computations primarily involve matrix multiplications (MM), which are efficiently handled by the inherent parallelism in PIM crossbar arrays where the activations are applied directly to the word-lines and summed along the bit-lines of crossbars [7]. Due to the wide range of devices supporting PIM technology, significant differences in the performance, specifically energy efficiency and latency overhead, are observed. Each device also has its shortcomings. For instance, SRAM consumes more power [9], whereas ReRAM suffers from higher latency and increased device noise [10]. To address these trade-offs, our approach leverages the strengths of both memories to achieve optimal performance.
II-B Photonic Tensor Cores and Modeling Challenges
Photonic tensor cores (PTCs) are emerging hardware platforms for ultra-fast matrix multiplication using light [11, 12, 13]. The diverse properties of PTC designs result in variations in circuit topology, devices, and operational principles, making accurate hardware modeling challenging. For example, based on expressivity, universal PTCs can map arbitrary matrix multiplication [11, 12], while subspace PTCs support only a subset of static linear transformation [14]. Based on the numerical range of input operands, full-range PTCs handle arbitrary values in one-shot computation, while subspace coherent PTCs require multiple computations to obtain the same results. Reconfiguration speed is another key factor to determine its supported dataflow and operations. Certain PTCs require thermal tuning to reconfigure the weights, leading to delays of and limiting them to weight-stationary dataflows, which are unsuitable for dynamic self-attention. Dynamic PTCs [15] use high-speed modulators for real-time matrix switching, enabling dynamic tensor products and output-stationary dataflows. For this work, we focus on using dynamic PTCs as a complement to weight-static electronic PIM.
II-C Prior Heterogeneous Accelerators
Several works proposed DNN accelerators that exploit heterogeneity in PIM devices and digital architectures (e.g., GPUs, TPUs) to optimize performance, efficiency, and accuracy.
Significant efforts have also been made to mitigate non-idealities in ReRAM-based PIM.
For example, some designs employ SRAM for most-significant-bit calculations and ReRAM for least-significant-bit calculations following weight quantization [16], improving accuracy at the cost of higher power consumption and increased area.
Meanwhile, research such as HyDe [17] proposes to optimize the design with a single-objective (e.g., latency or accuracy) by identifying the optimal layer-to-device mapping, though it struggles in non-convex search regions.
Other works [18] have proposed a heterogeneous architecture that utilizes PIM for weight-stationary operations and TPUs for dynamic matrix multiplication in transformer models.
To the best of our knowledge, no existing research integrates photonics, which offers superior overall performance compared to TPUs, with PIM devices for DNN acceleration.
These gaps in research motivate us to introduce H3PIMAP, a mapping framework for efficient partitioning and mapping of DNN workloads to heterogeneous electronic-photonic-PIM accelerators.
III Proposed H3PIMAP Framework on Electronic-Photonic-PIM Architecture
We first illustrate the heterogeneous architecture with feature and noise analysis and then introduce our proposed mapping framework H3PIMAP for efficient, robust hybrid DNN workload mapping.
III-A Architecture Overview and Analysis
As a case study, we consider a multi-tier 3D-stacked accelerator architecture with ReRAM PIM tier (R), SRAM PIM tier (S), integrated photonics computing tier (P), and global buffer (M) tier, shown in Fig. 1. Each tier contains a multi-tile accelerator with one type of dedicated technology (R/S/P). Tiers are interconnected via silicon through vias (TSVs) for high-bandwidth data movement. The hierarchical on-chip global buffer for weights (needed by photonics) and activations (needed by all three computing tiers) is located at the memory tier (M).
As an important motivation, we answer a critical question: why do we need heterogeneous architectures that integrate three technologies (R/S/P)?
Feature Analysis of 3 Technologies – Key features of accelerators based on three technologies are summarized in Table I, highlighting key trade-offs that motivate the need for a heterogeneous architecture with a heterogeneity-aware mapping framework.
@@ III-A Architecture Overview and Analysis
➊ Speed: Photonic accelerator exhibits unparalleled computational speed even with a few small-size tensor cores, achieving 110 TOPS performance. This makes them highly suitable for compute-bound workloads with high arithmetic intensity, e.g., convolution and self-attentions. In contrast, ReRAM-based PIM architectures typically show low operation speed due to their high latency and multiplexed ADC sampling mode, but excel in energy efficiency for memory-bound operations, i.e., large matrix multiplication operations, making them an ideal choice for efficiency-prioritized tasks. SRAM PIM accelerators fall between the other two technologies, offering balanced performance for moderately sized workloads.
➋ Weight-dynamic vs. Weight-static Operation Support: The inherent characteristics of these technologies dictate their suitability for different workload types. ReRAM and SRAM PIM architectures are well-suited for weight-static operations, where weights are mapped to dedicated PIM PEs to minimize frequent weight updates due to its weight-stationary dataflow and ReRAM endurance limits. On the other hand, dynamic operations, such as self-attention in Transformers, require frequent updates on both operands (e.g., Q and K tensors). are better handled by dynamic photonic PEs.
➌ Bit Resolution and Noise Robustness: SRAM and ReRAM PIM accelerators provide computation for 8-bit integer multiplication due to their high cell density and precise digital mechanisms. However, the 2-bit ReRAM cells are susceptible to thermal noises with conductance drift [10].
Both PIM architectures generate significant heat during operations, which can degrade the overall robustness of the PEs. This heat impacts not only the current tier where the computation is taking place but can also propagate to adjacent tiers in 3D-stacked architectures, compromising accuracy due to slow thermal dissipation. Photonic tensor cores, while capable of one-shot light-speed matrix multiplications, are typically limited to lower bit resolutions (e.g., <6-bit operands). Additionally, they are susceptible to hardware noise, which can affect computational accuracy.
III-B Dataflow and Interconnect Design
@@
To match the distinct properties and operational demands of both PIM and photonic accelerators, we customize a dataflow to partition matrix multiplication in each DNN layer among three computing tiers (R/S/P). As each PIM tile communicates solely with the global buffer via TSVs and routers, data transfer remains one-dimensional. This approach eliminates inter-tile communication overhead. To further optimize communication efficiency, TSV connections to the PIM tiers are positioned midway between PIM tiles. This placement effectively halves the average communication distance relative to a 2D network-on-chip. Additionally, we add a dedicated TSV link to the photonic tier, accommodating its substantial bandwidth requirements. Our following introduced mapping algorithm and system evaluations are based on the above dataflow and interconnect designs.
III-C Non-ideal Hardware Noise Modeling
As our architecture contains analog computing hardware (ReRAM and photonics), their hardware non-ideality, e.g., noise variation, will impact the computing fidelity and model inference accuracy. Their unique noise properties imply that a noise-aware mapping strategy can boost inference accuracy. To understand the hardware robustness, we utilize standard models to capture device-level noise variations across different tiers.
PIM: We consider SRAM to be robust to thermal noises [9] for the scope of this work due to its high thermal tolerance and digital computing mechanism. However, ReRAM is thermally sensitive. Major sources of noise in ReRAM [10] are thermal and shot noise, formulated as
@@ III-C Non-ideal Hardware Noise Modeling
(1) where , , , , and denote the conductance, terminal voltage, Boltzmann constant, temperature, and frequency, respectively.
@@
Photonics: We adopt the TeMPO architecture [8] and its real-device noise measurements as our noise model. Following its original settings, the photonic noise is treated as a relative random Gaussian perturbation injected into both the input tensors of matrix multiplications, formally modeled as , where . According to TeMPO, real measurements under typical operating conditions show that is most representative, and thus we adopt as photonic tier noise.
III-D Mapping Problem Formulation
@@
Given an -layer DNN with sequential layer execution, our goal is to determine an optimal matrix row-to-tier mapping individually for each layer on our heterogeneous architecture. Each weight matrix row is assigned to one of the computing tiers. We use to represent the percentage of matrix rows mapped to tier . The objective of the mapping process is to minimize the overall model inference energy () and latency () under several specific constraints.
This can be formulated as the following multi-objective optimization (MOO) problem:
@@
The final configuration specifies how each layer’s weight rows are assigned across tiers. The total inference latency equals the sum of individual layer latencies, with each layer bottlenecked by the slowest tier. The overall energy consumption is computed by adding the energy usage of all layers. Here, denotes the total memory capacity (or tile size) of tier . To avoid storage overflow, no tier may be assigned more weight rows than it can accommodate. As discussed in ➋, any operation that a given tier cannot support will not be mapped to that tier. Because device noise and limited precision can degrade performance, we require to remain within discrepancy below the original accuracy .
@@
Search Space Analysis: The complexity of the problem can be illustrated with the following example. For an -layer DNN with neurons (weight rows) per layer mapped onto tiers, there are total possible mappings. For example, a popular language model Pythia-70M has an average of 2048 neurons per layer and 6 layers in total. With 3 tiers to map, the total search space is 312288. Besides the vast search space, the non-trivial cost of evaluating the accuracy and hardware efficiency of each solution cast significant challenges to the MOO task. To efficiently explore the design space of this constrained multi-objective optimization problem, we propose H3PIMAP, a two-stage search and optimization procedure as shown in Fig. 2. We decouple the accuracy and efficiency optimization in two stages and prune the search space with sensitivity-aware heuristics. The first stage rapidly explores the energy-latency space to obtain Pareto-optimal mapping candidates. Then, the second stage performs accuracy-driven row remapping to strategically calibrate the row assignment to satisfy the accuracy constraints.
III-D1 Stage 1: Latency-Energy Pareto Optimization (PO)

### v2

# HPIMAP: A Heterogeneity-Aware Multi-Objective DNN Mapping Framework on Electronic-Photonic Processing-in-Memory Architectures

@@ Abstract.
The future of artificial intelligence (AI) acceleration demands a paradigm shift beyond the limitations of purely electronic or photonic architectures. Photonic analog computing delivers unmatched speed and parallelism but struggles with data movement, robustness, and precision. Electronic processing-in-memory (PIM) enables energy-efficient computing by co-locating storage and computation but suffers from endurance and reconfiguration constraints, limiting it to static weight mapping. Neither approach alone achieves the balance needed for adaptive, efficient AI. To break this impasse, we study a hybrid electronic-photonic-PIM computing architecture and introduce H3PIMAP, a heterogeneity-aware mapping framework that seamlessly orchestrates workloads across electronic and optical tiers. By optimizing workload partitioning through a two-stage multi-objective exploration method, H3PIMAP harnesses light speed for high-throughput operations and PIM efficiency for memory-bound tasks. In system-level evaluations, H3PIMAP delivers a 3.32× latency reduction across language and vision models and, on large language models, achieves 77.0 lower latency with 14.6 lower energy at matched quality, outperforming homogeneous and naïve mappings strategies. This proposed framework lays the foundation for hybrid AI accelerators, bridging the gap between electronic and photonic computation for next-generation efficiency and scalability.
1. Introduction
@@
Artificial Intelligence (AI), powered by deep neural networks (DNNs), now underpins a vast spectrum of applications, from computer vision and natural language processing (NLP) to medical diagnostics and time-series analytics. These workloads increasingly demand sub-millisecond latency, high throughput, and stringent energy efficiency, requirements that traditional CPU/GPU-based architectures often struggle to satisfy. The growing disparity between processing speed and data-movement cost, the notorious memory wall, has motivated a paradigm shift toward domain-specific accelerators that transcend the classical von Neumann model.
Among these emerging paradigms, Processing-in-Memory (PIM) and photonic accelerators have attracted particular attention for their complementary strengths. PIM architectures exploit the intrinsic capability of performing arithmetic directly within or near memory arrays, drastically reducing costly data transfers between separate processing and memory units. By merging computation and storage, PIM architectures can significantly cut both latency and energy consumption (Sebastian et al., 2020). However, the landscape of PIM technologies, ranging from SRAM (Zhang et al., 2017) to FeFET (Long et al., 2019), ReRAM (Feinberg et al., 2018), and MRAM (Yusuf et al., 2024), presents intricate trade-offs in area, endurance, write latency, and power dissipation. No single technology dominates across all metrics, underscoring the need for careful technology-architectural co-design to tailor PIM solutions to specific workload demands.
Photonic accelerator systems exploit light‐based signal propagation and processing to target the heavy linear algebra that dominates DNN inference and training. For example, by using integrated photonic circuits to carry out matrix multiplication and accumulation, photonic computing has demonstrated sub-nanosecond inference latency and very low energy per operation. Yet, translating these proof-of-concept results into large-scale, general-purpose DNN accelerators remains challenging. Key challenges include fabrication variation and non-ideal noise in optical modulators, thermal drift and crosstalk across components, limited numerical precision on high-speed optical signal modulation, and the overhead of electro-optic signal conversion. Overcoming these limitations requires robust calibration, error-tolerance, and co-design strategies to fully unlock photonics’ potential in AI acceleration. The complementary capabilities of photonics and electrical PIM offer an opportunity to leverage the best of both worlds to explore a hybrid computing paradigm.
On the application level, the inherent heterogeneity of modern DNNs presents both challenges and opportunities of co-design. Different model layers and modules exhibit varying levels of data reuse, computational intensity, and sensitivity. For instance, mapping highly dynamic attention layers in Transformers onto non-volatile PIM arrays can lead to rapid device wear-out (Chi et al., 2016), while photonics might excel for dense, static matrix multiplications but is less suited for depthwise or data-transformation-heavy operations. These diverse workload characteristics of modern DNN models imply a perfect match with heterogeneous electronic-photonic-PIM hardware architecture, but demand heterogeneity-aware mapping strategies that intelligently allocate tasks to the most suitable accelerator type, maximizing efficiency without compromising accuracy or device lifetime.
To enable efficient workload mapping onto next-generation electronic-photonic-PIM accelerators, we propose H3PIMAP, a heterogeneity-aware multi-objective mapping scheme, an effective two-stage mapping flow that jointly optimizes energy, latency, and robustness:
1. Pareto-optimal exploration stage identifies optimal energy-latency configurations across hybrid PIM-photonic platforms using integrated architectural and interconnect simulation.
2. Sensitivity-aware remapping stage refines the initial mapping to preserve model accuracy under hardware non-idealities, ensuring resilience against device variability and quantization effects.
Through this integrated design and mapping approach, H3PIMAP enables efficient cooperation between electronic and photonic accelerators for hybrid AI workloads. Our key contributions are as follows:
@@
Electronic-Photonic-PIM Accelerator Evaluation Infrastructure: We create the first system modeling and optimization framework integrating SoTA simulators for PIM, photonic AI hardware,
and network-on-chips, enabling automated design exploration.
@@
Heterogeneity-aware Mapping Framework: We introduce a two-stage workload mapping framework that considers a workload’s variability and target hardware’s heterogeneity. This ensures rapid mapping for space exploration for optimal resource utilization, system performance, and inference accuracy.
Comprehensive Performance Evaluation: Across diverse NLP and vision benchmarks, H3PIMAP delivers 3.32 lower latency than homogeneous systems and naïve mappings; on large language models (LLMs), the gains are even stronger, 77.0 lower latency and 14.6 lower energy under matched-quality constraints.
2. Related Work
We first survey DNN accelerators built entirely from a single technology—SRAM-, ReRAM-based PIM, or photonics, and compare their characteristics in Table 1.
PIM Accelerators: Numerous PIM (Zhang et al., 2017; Long et al., 2019; Feinberg et al., 2018; Yusuf et al., 2024; Shafiee et al., 2016; Chi et al., 2016; Shaik et al., 2024) designs push matrix multiplication directly into memory arrays, cutting data movement but trading off write latency, endurance, or noise. Recent surveys (Joardar et al., 2020) indicate that no single device dominates all metrics, motivating the adoption of hybrid solutions.
Photonic Tensor Cores: Photonic tensor cores (Xu et al., 2021; Shen et al., 2017; Tait et al., 2017; Feng et al., 2022; Zhu et al., 2024) deliver sub-nanosecond reconfiguration, passive device matrix computation, and >800 TOPS throughput. Their combined power-performance-area (PPA), therefore, surpasses any known digital accelerator. Nevertheless, thermal drift and ADC/DAC overheads still cap full-system efficiency, and faithfully modeling these non-idealities remains an open challenge.
Hybrid & Heterogeneous Accelerators: Prior work couples SRAM with ReRAM for MSB/LSB splitting or combines GPUs with PIM arrays, yet largely optimizes a single objective or ignores photonics altogether. (Rashed et al., 2021; Bhattacharjee et al., 2023; Luo and Yu, 2024) To the best of our knowledge, no prior work integrates photonics with PIM devices for DNN acceleration while simultaneously tackling layer-to-device assignment under joint latency, energy, and accuracy constraints, a gap that motivates our H3PIMAP framework for efficiently partitioning and mapping DNN workloads onto heterogeneous electronic–photonic–PIM accelerators.
3. Proposed H3PIMAP Framework on Electronic-Photonic-PIM Architecture
In this section, we present a heterogeneous electronic-photonic-PIM architecture, analyze its key characteristics and device noise, and describe H3PIMAP framework that exploits these features for efficient, robust DNN mapping.
3.1. Architecture Overview and Analysis
As a concrete case, we consider a four-tier 3D-stacked accelerator architecture: a ReRAM-PIM layer (R), an SRAM-PIM layer (S), an integrated photonic layer (P), and a global-buffer layer (M). Each tier hosts many identical tiles built in a single technology and silicon through vias (TSVs) ferry data between tiers. The shared buffer in tier M holds photonic weights and activations for all three compute layers.
As an important motivation, we answer a critical question: Why bother fusing R, S, and P into a heterogeneous architecture?
Feature Analysis of 3 Technologies – Table 1 highlights the three technologies’ speed, energy, and precision in complementary ways, compelling a heterogeneity-aware mapper.
@@ 3.1. Architecture Overview and Analysis
➊ Speed: Photonic cores race through math, 1–10 TOPS with only a handful of tensor units, so they shine on compute-bound layers (convs, self-attention). ReRAM PIM is slower, throttled by ADC sampling, yet fantastically energy-wise for memory-bound matrix multiplies. SRAM PIM lands between, giving balanced throughput for medium workloads.
➋ Static vs. dynamic weights: ReRAM and SRAM favor weight-static workloads, weights stay resident to keep device endurance (e.g., ReRAM). Dynamic layers that rewrite both operands (e.g., Q K in Transformers) suit photonic cores better.
➋ Precision and noise: SRAM/ReRAM handle 8-bit MACs but run hot, and 2-bit ReRAM cells drift. Photonic cores perform one-shot light-speed multiplies yet are limited to <6 bits and suffer thermal crosstalk and fabrication noise—calibration is essential.
3.2. Dataflow and Interconnect Design
@@
To match the distinct properties and operational demands of both PIM and photonics, we customize a dataflow to partition matrix multiplication in each DNN layer among three computing tiers (R/S/P). As each PIM tile communicates solely with the global buffer via TSVs and routers, data transfer remains one-dimensional. This approach eliminates inter-tile communication overhead. To further optimize communication efficiency, TSV connections to the PIM tiers are positioned midway between PIM tiles. This placement effectively halves the average communication distance relative to a 2D network-on-chip. Additionally, we add a dedicated TSV link to the photonic tier, accommodating its substantial bandwidth requirements. Our following introduced mapping algorithm and system evaluations are based on the above dataflow and interconnect designs.
3.3. Non-ideal Hardware Noise Modeling
Analog ReRAM and photonic tiers suffer thermal, shot, and phase noise that threaten accuracy, so we apply established per-tier noise models to steer a robustness-aware mapper.
PIM: SRAM’s high tolerance and digital logic render it immune to thermal noise (Shaik et al., 2024), while thermally sensitive ReRAM faces dominant thermal and shot noise (Joardar et al., 2020), formulated as:
@@ 3.3. Non-ideal Hardware Noise Modeling
(1) where , , , , and are conductance, voltage, Boltzmann constant, temperature, and frequency, respectively.
@@
Photonics: To model photonic noise, we adopt the TeMPO architecture (Zhang et al., 2024) and its device-level measurements. Following TeMPO, we model noise as a relative Gaussian perturbation on each matrix input, i.e., , where . We fix for the photonic tier.
3.4. Mapping Problem Formulation
@@
Given an -layer DNN with sequential layer execution, our goal is to obtain an optimal matrix row-to-tier mapping individually for each layer on our heterogeneous architecture. Each matrix row is assigned to one computing tier. We use to denote the percentage of matrix rows mapped to tier . The objective of mapping is to minimize the overall model inference energy () and latency () under several specific constraints. This can be formulated as the following multi-objective optimization (MOO) problem:
@@
The final configuration specifies how each layer’s weight rows are assigned across tiers. Total latency is the sum of layer latencies, each gated by its slowest tier. Total energy is the sum of per-layer energies. As discussed in Section 3.1➋, any operation that a given tier cannot support will not be mapped to that tier. Because device noise and limited precision can degrade task performance, our noise-aware mapping framework requires the post-mapping accuracy to remain within discrepancy below the original pretrained accuracy .
@@
Search Space Analysis: The complexity of the problem can be illustrated with the following example. For an -layer DNN with neurons (weight rows) per layer mapped onto tiers, there are total possible mappings. For example, a popular language model Pythia-70M has an average of 2048 neurons per layer and 6 layers in total. With 3 tiers to map, the total search space is 312288. Besides the vast search space, the non-trivial cost of evaluating the accuracy and hardware efficiency of each solution poses significant challenges to the MOO task. To efficiently explore the design space of this constrained multi-objective optimization problem, we propose H3PIMAP, a two-stage search and optimization procedure as shown in Fig. 2. We decouple the accuracy and efficiency optimization in two stages and prune the search space with sensitivity-aware heuristics. The first stage rapidly explores the energy-latency space to obtain Pareto-optimal mapping candidates. Then, the second stage performs accuracy-driven row remapping to strategically calibrate the row assignment to satisfy the accuracy constraints.
3.4.1. Stage 1: Latency-Energy Pareto Optimization (PO)